### v2

# Strategic Data Center Load Shifting: Implications for Market Efficiency and Transmission Value

Abstract
Data center electricity use may reach 12% of U.S. demand by 2030, alongside growing ability to shift workloads geographically in response to prices or carbon signals. We examine the system-level implications of such strategic flexibility using a bilevel two-zone model that couples economic dispatch with consumer cost minimization. Two market failures emerge. First, discontinuous price changes at generator capacity limits can induce flexible consumers to shift load in socially inefficient directions—for example, toward a higher-cost region to trigger a price drop elsewhere. Second, by positioning near capacity boundaries, consumers can eliminate the marginal value of transmission expansion: although shadow prices suggest additional capacity is valuable, consumers reoptimize to offset resulting flow changes, leaving dispatch and costs unchanged. We derive conditions under which these effects arise and show that conventional price signals and planning methods can misrepresent system value in the presence of large flexible loads.
I Introduction
Data center growth is rapidly reshaping electricity demand. In the U.S., data center consumption may more than triple from 2023 to 2030, reaching $\sim$12% of total demand [mckinsey2024]. Because large operators can migrate workloads, these facilities act not only as major loads but also as sources of spatial and temporal flexibility—capabilities already used to reduce costs or emissions of hyperscale computing [google_env_report, Meta_sust_report, ZhengEtal2020_joule, radovanovic2022carbon].
A common assumption is that this flexibility should improve system efficiency: by shifting load toward locations with lower locational marginal prices (LMPs), flexible consumers will provide a similar benefit to the system as transmission—reducing congestion and better utilizing cheap generation [zhang2020flexibility]. This view, however, assumes consumers are price-takers whose decisions do not influence dispatch. As data center loads grow large enough to affect which generators are marginal, decentralized shifting can alter prices and congestion in ways that may not align with system objectives.
These observations motivate two central questions:
  1. 1.
    When can decentralized, price-responsive load shifting increase system costs?
  2. 2.
    How do such shifting incentives reshape the effective value of transmission?
The second question follows from the first: if flexible load can change dispatch outcomes, the benefits of transmission expansion may depend on consumers’ strategic responses.
Prior work exploring benefits of flexibility often assumes centralized coordination [zhang2020flexibility, knittel2025] or treats consumers as price/emissions-takers [radovanovic2022carbon, SouzaEtal2023], and consequently does not capture such strategic interactions. While recent work identifies counterintuitive outcomes from decentralized shifting [GorkaEtal2024], the mechanisms by which flexible consumers exert market power, and their implications for transmission value, remain underexplored.
We address this gap using a bilevel model in which a flexible consumer chooses spatial load allocation while the system operator performs economic dispatch. In a two-zone system, we derive closed-form conditions under which decentralized incentives align with or diverge from system objectives. Our contributions are:
1. Strategic misalignment. We show that decentralized shifting can raise system costs even as flexible consumers lower their own. Because consumers pay marginal-cost-based prices for load while the system incurs marginal costs only for incremental dispatch, consumers can benefit from positioning near generator capacity limits to depress prices (Theorem 1).
2. Zero bilevel marginal value of transmission. At these strategic equilibria, incremental transmission expansion has zero marginal value: flexible consumers shift load one-for-one with new line capacity, offsetting its effect and leaving dispatch unchanged, even as shadow prices remain positive.
3. Implications for planning and design. Because marginal pricing can induce strategic behavior that departs from socially optimal outcomes and biases transmission investment signals, our findings point to a potential need to refine planning and market models to better account for strategic flexible demand.
II Two-Zone Bilevel Model
We model the interaction between a flexible electricity consumer and a system operator as a bilevel optimization problem. The consumer (upper level) chooses how to allocate load across two zones to minimize procurement costs, while the system operator (lower level) performs economic dispatch given the consumer’s allocation decision.
Generation and network structure. Consider a power system with two zones, $A$ and $B$, each containing a merit-order supply stack of generation resources. Let $\mathcal{G}_{A}$ and $\mathcal{G}_{B}$ denote the ordered sets (with respect to marginal cost) of generators at zones $A$ and $B$ respectively. Each generator $g\in\mathcal{G}_{A}\cup\mathcal{G}_{B}$ is characterized by: (i) the capacity constraint: $0\leq p_{g}\leq P_{g}$ (MW), and (ii) constant marginal generation costs: $C_{g}\geq 0$ (\$/MWh). Let $F\geq 0$ denote the total transmission capacity between the two zones, and let $f\in[-F,F]$ denote the power flow from zone $A$ to zone $B$, where positive values indicate AA-to-$B$ flow.
Load allocation. Each zone has two types of loads: (i) Inflexible base load: $b_{A},\,b_{B}\geq 0$ (MWh), which cannot be shifted; (ii) Flexible consumer load: Initially allocated as $x_{A},\,x_{B}\geq 0$ (MWh). Importantly, the flexible consumer is able to shift load between zones. Let $\delta\in[-\alpha,\alpha]$ denote the net load shift from zone $A$ to zone $B$, where $\delta>0$ (resp. $\delta<0$) means net shift from $A$ to $B$ (resp. from $B$ to $A$). Here, $\alpha>0$ denotes the maximum shifting capability. After shifting, the realized flexible loads are $x_{A}-\delta$ at zone $A$ and $x_{B}+\delta$ at zone $B$.
II-A Lower Level: Economic Dispatch
Given a load shift $\delta$, the system operator solves an economic dispatch to minimize total generation cost:
Definition 1 (Market Clearing Problem).
$G(\delta,F):=\min_{\boldsymbol{p},\,f}\quad$ $\sum_{g\in\mathcal{G}_{A}\cup\mathcal{G}_{B}}C_{g}\,p_{g}$ (1a)
$\mathrm{s.t.}\quad$ $\sum_{g\in\mathcal{G}_{A}}p_{g}=b_{A}+x_{A}-\delta+f\quad(\lambda_{A})$ (1b)
$\sum_{g\in\mathcal{G}_{B}}p_{g}=b_{B}+x_{B}+\delta-f\quad(\lambda_{B})$ (1c)
$-F\leq f\leq F\hskip 51.21504pt(\mu^{-},\mu^{+})$ (1d)
$0\leq p_{g}\leq P_{g},\hskip 31.2982pt\forall g\in\mathcal{G}_{A}\cup\mathcal{G}_{B}$ (1e)
The dual variables have standard economic interpretations: $\lambda_{A},\lambda_{B}$ denotes the Locational Marginal Prices (LMPs) at zones $A$ and $B$; $\mu^{+},\mu^{-}$ the shadow price of upper/lower transmission limits. Importantly, LMPs equal the marginal cost of the marginal generator in each zone.
II-B Upper Level: Flexible Consumer Problem
The flexible consumer chooses a load shift $\delta$ to minimize total procurement cost:
Definition 2 (Consumer Cost Minimization).
Let $\lambda_{A}(\delta,F)$ and $\lambda_{B}(\delta,F)$ denote the LMPs resulting from dispatch problem (1). The consumer’s cost is:
$P(\delta,F):=\lambda_{A}(\delta,F)(x_{A}-\delta)+\lambda_{B}(\delta,F)(x_{B}+\delta)$ (2)
The consumer solves:
$\min_{-\alpha\leq\delta\leq\alpha}\quad P(\delta,F)$ (3)
Unlike the system objective $G(\delta,F)$, which is convex and continuous in $\delta$, the consumer objective $P(\delta,F)$ exhibits two key properties:
  1. 1.
    Nonconvexity: $P(\delta,F)$ is generally nonconvex due to the dependence on endogenous prices $\lambda_{i}(\delta,F)$
  2. 2.
    Discontinuities: When $\delta$ crosses a threshold where the marginal generator changes, the LMPs jump discontinuously. This causes $P(\delta,F)$ to have upward (resp. downward) jumps at these thresholds where more (resp. less) expensive generators become marginal.
These discontinuities are the fundamental source of misalignment between private and social incentives, which we analyze in Sec. III.
II-C Objective Alignment
Definition 3 (Objective Alignment).
Fix a reference point $\delta_{0}$ (e.g., $\delta_{0}=0$ for no shifting). Define the incremental changes:
$\Delta P(\delta)$ $:=P(\delta,F)-P(\delta_{0},F)$ (4)
$\Delta G(\delta)$ $:=G(\delta,F)-G(\delta_{0},F)$ (5)
We say the system and consumer objectives are aligned with respect to shift $\delta$ if:
$\Delta P(\delta)<0\quad\Rightarrow\quad\Delta G(\delta)\leq 0,$ (6)
and misaligned otherwise (i.e., if the flexible consumer prefers a shift that increases system cost).
While deliberately stylized, our two-zone framework captures the key mechanisms of price formation and congestion in power systems while remaining analytically tractable. It also reflects two practically relevant settings for hyperscale load shifting. First, setting $F=0$ models two independent electricity markets (e.g., different continents or asynchronous grids) with no physical power transfer, isolating how differences in local supply stacks—particularly the slope of merit order curves—can create misalignment between system and consumer objectives. Second, setting $F>0$ represents two regions within an interconnected market, highlighting how spatially flexible demand interacts with congestion and marginal pricing to distort the perceived value of transmission capacity and alter investment signals.
III Disconnected Electricity Markets
In the following section, we study the case of two disconnected electricity markets (i.e., assuming $F=0$) and identify conditions under which system generation cost and flexible consumer cost objectives are aligned or misaligned. For now, we adopt the notation $G(\delta)\coloneqq G(\delta,0)$ and $P(\delta)\coloneqq P(\delta,0)$.
III-A Basis Changes and Discontinuities
When load shifts are large enough to change marginal generators, the flexible consumer cost function develops discontinuities that yield potential for misalignment. In Proposition 1, we describe how system and flexible consumer costs change when flexible consumers exert market power over prices. Specifically, Proposition 1 relates changes in the system and flexible consumer objectives at basis changes of (1), denoted by $\tau_{1},\,\tau_{2}$.
Proposition 1.
Consider a positive load shift $\delta>0$ from $A$ to $B$. Letting $p_{g_{A}},\,p_{g_{B}}$ respectively denote the pre-shift dispatch levels of marginal generators $g_{A}$ and $g_{B}$, two cases arise depending on which zone’s marginal generator changes first.
Case 1. (Marginal generator at $B$ changes first). Suppose
$P_{g_{B}}-p_{g_{B}}<p_{g_{A}}.$
Let $\tau_{1}:=P_{g_{B}}-p_{g_{B}}$ be the threshold where generator $g_{B}$ reaches capacity. Then, for $\delta\leq\tau_{1}$:
$\Delta G(\delta)=\begin{cases}\Delta P(\delta),&\delta\in[0,\tau_{1}),\\[2.0pt] \Delta P(\delta)-\kappa_{1},&\delta=\tau_{1},\end{cases}$
where the discontinuity $\kappa_{1}$ is given by
$\kappa_{1}\coloneqq(C_{g_{B+1}}-C_{g_{B}})(x_{B}+\tau_{1})>0.$
Case 2. (Marginal generator at $A$ changes first). Suppose
$p_{g_{A}}<P_{g_{B}}-p_{g_{B}}.$
Let $\tau_{2}:=p_{g_{A}}$ be the threshold where generator $g_{A}$ is fully de-committed. Then, for $\delta\leq\tau_{2}$:
$\Delta G(\delta)=\begin{cases}\Delta P(\delta),&\delta\in[0,\tau_{2}),\\[2.0pt] \Delta P(\delta)+\kappa_{2},&\delta=\tau_{2},\end{cases}$
where the discontinuity $\kappa_{2}$ is given by
$\kappa_{2}\coloneqq(C_{g_{A}}-C_{g_{A-1}})(x_{A}-\tau_{2})>0.$
Interpreting Proposition 1, we first note that, for $\delta$ smaller than the first basis-change threshold (i.e., before either $\tau_{1}$ or $\tau_{2}$ is reached):
$\Delta G(\delta)=\Delta P(\delta)=(C_{g_{B}}-C_{g_{A}})\,\delta.$
In other words, when load shifts are small enough that the optimal basis of the dispatch problem does not change, system and consumer objectives are perfectly aligned: both benefit linearly from shifting toward the lower-cost region. In this regime, marginal generation costs act as fully welfare-aligned price signals.
Once a basis change occurs, however, Proposition 1 reveals a fundamental asymmetry. The system pays the cost of a newly marginal generator only on its incremental production, whereas the consumer pays the new marginal price on all load at that location. This discrepancy creates discontinuous jumps in the consumer objective at $\delta=\tau_{1}$ or $\delta=\tau_{2}$. As a consequence, flexible consumers may find it profitable to depress prices by strategically shifting load to alter which generators are marginal in zones $A$ and $B$, even when such shifts increase total system cost.
III-B Strategic Positioning and Misalignment
Theorem 1 formalizes the conditions under which the discontinuities identified in Proposition 1 lead to negative externalities, causing decentralized load shifting to diverge from system-optimal behavior.
Theorem 1 (Externalities at basis changes).
Consider a positive load shift $\delta=\tau>0$ from $A$ to $B$ that is exactly the minimal shift required to change the optimal basis of (1) (i.e., to induce a change in the marginal generator at either $A$ or $B$). Let $p_{g_{A}}$ and $p_{g_{B}}$ denote the pre-shift dispatch levels of the marginal generators at $A$ and $B$.
The system and consumer objectives are aligned for all such shifts except when all of the following conditions hold:
  1. (i)
    The minimal basis-changing shift causes de-commitment of the marginal generator at $A$:
    $\tau=p_{g_{A}}<P_{g_{B}}-p_{g_{B}}.$
  2. (ii)
    Before the shift, the marginal generator at $B$ is strictly more expensive:
    $C_{g_{B}}>C_{g_{A}}.$
  3. (iii)
    The downward jump in the price at $A$ (from switching $g_{A}$ to $g_{A-1}$) offsets the increase in price exposure at $B$:
    $(C_{g_{A}}-C_{g_{A-1}})(x_{A}-\tau)>(C_{g_{B}}-C_{g_{A}})\,\tau.$
When these three conditions hold, the flexible consumer reduces their procurement cost at $\delta=\tau$ while system generation cost increases, yielding misalignment.
Theorem 1 can be interpreted as follows. A flexible consumer may be incentivized to shift load against the direction suggested by the LMP difference when one or more of the following conditions hold:
  1. 1.
    Large cost drop at $\boldsymbol{A}$: the difference $(C_{g_{A}}-C_{g_{A-1}})$ is large.
  2. 2.
    Small inter-zone price difference: the gap $(C_{g_{B}}-C_{g_{A}})$ is small.
  3. 3.
    Large allocation at $\boldsymbol{A}$: the quantity $(x_{A}-\tau_{2})$ is large, so the resulting price drop applies to many megawatt-hours.
  4. 4.
    Small shift required: the threshold $\tau_{2}$ is small, making it easy to trigger the marginal generator change.
Under these conditions, local marginal prices fail to encode system-optimal incentives, and decentralized load shifting can increase total system cost.
III-C Illustrative Numerical Example
We illustrate this phenomenon in Fig. 1 using a simple two-zone example in which flexible load accounts for 20% of total demand. The system parameters are:
Base loads: $b_{A}=b_{B}=800$ MWh
Flexible loads: $x_{A}=400$ MWh and $x_{B}=0$ MWh
Zone $\boldsymbol{A}$ generators:
  1. 1.
    $g_{A}^{1}$: 1 GW at \$0/MWh
  2. 2.
    $g_{A}^{2}$: 200 MW at \$25/MWh
Zone $\boldsymbol{B}$ generators:
  1. 1.
    $g_{B}^{1}$: 1.2 GW at \$40/MWh
In this configuration, a flexible consumer minimizes procurement cost by shifting half of their load from $A$ to $B$ so that generator $g_{A}^{2}$ is no longer marginal. This reduces the LMP at $A$ from \$25/MWh to \$0/MWh. However, the resulting load shift forces a larger share of energy to be supplied by $g_{B}^{1}$ at \$40/MWh, increasing total system generation costs. This occurs because the cost gap between $g_{A}^{1}$ and $g_{A}^{2}$ is larger than the gap between $g_{A}^{2}$ and $g_{B}^{1}$, creating an incentive for strategic shifting even though it raises system costs.
Figure 1: Comparison of system generation costs (left) and flexible consumer costs (right) for a range of shifts $\delta\in[0,250]$.
IV Transmission-Coupled Markets
We now introduce transmission coupling ($F>0$), which creates additional strategic interactions beyond those in Section III. When load is inflexible, the short-run marginal value of transmission—defined by $\frac{d}{dF}G$, or the reduction in cost per unit of added capacity—is conventionally understood to be the shadow price of the binding transmission constraint [hogan1992contract]. With flexible load, however, locational prices influence the spatial allocation of demand, meaning that transmission and load shifting interact. This raises the question of how price-responsive load shifting affects the marginal value of increasing transmission capacity.
IV-A Marginal Value of Transmission Under Flexible Load
Theorem 2 addresses this question in a regime where flexible load optimally positions at a basis change by de-committing the otherwise-marginal generator at $A$ (see Theorem 1). Under these conditions, transmission upgrades no longer yield the cost reductions suggested by the shadow price.
Theorem 2 (Zero bilevel marginal value of transmission).
Fix initial transmission capacity $F_{0}\geq 0$ and let $\delta_{0}:=\delta^{*}(F_{0})$ be the consumer-optimal load shift. Suppose the following conditions hold:
  1. (A1)
    Marginal generation costs satisfy $C_{g_{A}}<C_{g_{A+1}}<C_{g_{B}}$ and the transmission line is congested from $A$ to $B$:
    $\lambda_{A}=C_{g_{A}}<C_{g_{B}}=\lambda_{B},\qquad f=F_{0}.$
  2. (A2)
    Strategic positioning is profitable:
    $(C_{g_{A+1}}-C_{g_{A}})\,(x_{A}-\delta_{0})>(C_{g_{B}}-C_{g_{A+1}})\,\delta_{0},$
    so that $g_{A}$ is being dispatched at level $p_{g_{A}}=P_{g_{A}}$.
  3. (A3)
    The marginal generator $g_{B}$ has downward headroom: $p_{g_{B}}>0$.
Define the threshold transmission capacity:
$\Gamma:=\left(\frac{C_{g_{A+1}}-C_{g_{A}}}{C_{g_{B}}-C_{g_{A}}}\right)\,x_{A}-\delta_{0}$
and the upper limit:
$F_{1}:=F_{0}+\min\left\{\Gamma,\,\alpha-\delta_{0}\right\}$
Then for all $F\in[F_{0},F_{1})$:
  1. (i)
    The transmission constraint remains binding with positive shadow price:
    $\mu^{+}=C_{g_{B}}-C_{g_{A}}>0.$
  2. (ii)
    The marginal value of transmission with respect to the system objective is zero under load shifting:
    $-\frac{d}{dF}\,G(\delta^{*}(F),F)=0.$
  3. (iii)
    The marginal value of transmission with respect to the flexible consumer objective is negative under load shifting:
    $-\frac{d}{dF}\,P(\delta^{*}(F),F)=C_{g_{A}}-C_{g_{B}}<0.$
In this regime, the shadow price $\mu^{+}$ no longer reflects the marginal value of transmission capacity. If all load were held fixed, increasing $F$ would allow zone $B$ to replace higher-cost local generation with lower-cost supply from zone $A$, yielding a reduction in system cost. However, at the equilibrium described in Theorem 2, the flexible load best response adjusts to keep generator $g_{A}$ marginal in zone $A$. As $F$ increases, the optimal load shift $\delta^{*}(F)$ increases one-for-one, leaving the effective net transfer between zones unchanged.
As a consequence, the dispatch and total system cost remain constant even as transmission capacity expands: the system-level marginal value of transmission is zero. Meanwhile, flexible consumer costs increase, since maintaining $g_{A}$ as the marginal generator in zone $A$ requires allocating more load to the higher-cost zone $B$. Thus, although transmission expansion appears beneficial when evaluated through its shadow price, this apparent benefit is fully offset by endogenous load shifting.
IV-B Illustrative Numerical Example
To illustrate the interaction between load shifting and transmission capacity, we consider a system with initial base and flexible loads identical to those in Sec. III-C and with the following generation resources.
Zone $\boldsymbol{A}$ generation resources:
  1. 1.
    $g_{1}^{A}$: 400 MW capacity at \$0/MWh
  2. 2.
    $g_{2}^{A}$: 600 MW capacity at \$25/MWh
  3. 3.
    $g_{3}^{A}$: Unlimited capacity at \$60/MWh
Zone $\boldsymbol{B}$ generation resources:
  1. 1.
    $g_{1}^{B}$: 200 MW capacity at \$0/MWh
  2. 2.
    $g_{2}^{B}$: 400 MW capacity at \$40/MWh
  3. 3.
    $g_{3}^{B}$: Unlimited capacity at \$70/MWh
We solve the market-clearing problem (1) over $\delta\in[0,400]$ and $F\in[0,200]$, and plot the resulting objective values $G(\delta,F)$ and $P(\delta,F)$ as heatmaps (Fig. 2). System operating conditions and cost metrics appear in Table I.
Across all $F$, the system-optimal load shift is $\delta=0$, which allows full use of cheaper generation at $A$ and avoids dispatching the high-cost unit $g_{B}^{3}$. However, when $F<\Gamma$ (107 MW), flexible consumers can lower their own costs by shifting load to $B$ to avoid dispatching $g_{A}^{3}$, thereby raising the price at $A$ from \$25/MWh to \$60/MWh. This blocks the system operator from serving $B$ with cheaper power from $g_{A}^{3}$, increasing total system cost (see $F=0$ and $F=100$ in Table I).
This illustrates the misalignment in objectives described in Theorem 2: the cost gap between marginal units within zone $A$ (between $g_{A}^{3}$ and $g_{A}^{2}$) is large relative to cross-zone differences, giving flexible load an incentive to shift even though it raises system cost. For $F\geq\Gamma$, this incentive disappears because serving load in zone $B$ becomes more expensive than depressing prices in zone $A$, restoring alignment between system and consumer objectives (see $F=200$).
Figure 2: System costs (left) and flexible consumer costs (right) evaluated over a range of load shifts $\delta$ and transmission capacities $F$.
Transmission Cap. (MW)
Metric $F=0$ $F=100$ $F=200$
No Load Shifting ($\delta=0$)
System Cost (\$) 57,000 56,000 55,000
Flexible Consumer Cost (\$) 24,000 24,000 24,000
LMPs $(\lambda_{A},\lambda_{B})$ (\$/MWh) (60, 70) (60, 70) (60, 70)
Marginal Units $g_{A}^{3},\,g_{B}^{3}$ $g_{A}^{3},\,g_{B}^{3}$ $g_{A}^{3},\,g_{B}^{3}$
Optimal Load Shifting
Optimal $\delta^{*}$ (MW) 200 300 0
System Cost (\$) 59,000 59,000 55,000
Flexible Consumer Cost (\$) 19,000 23,500 24,000
LMPs $(\lambda_{A},\lambda_{B})$ (\$/MWh) (25, 70) (25, 70) (60, 70)
Marginal Units $g_{A}^{2},\,g_{B}^{3}$ $g_{A}^{2},\,g_{B}^{3}$ $g_{A}^{3},\,g_{B}^{3}$
Impact of Load Shifting
$\Delta$ System Cost (\$) +2,000 +3,000 0
$\Delta$ Flexible Cost (\$) 5,000-5{,}000 500-500 0
TABLE I: Dispatch decisions and cost metrics for $F=0,100,200$.
The implications of Theorem 2 are further illustrated in Fig. 3, which compares system and flexible consumer costs as functions of $F$ both with and without load shifting. For $F<\Gamma$, endogenous load shifting exactly offsets the benefits of transmission expansion, yielding no reduction in system cost while raising costs for the flexible consumer. For $F\geq\Gamma$, this offset vanishes and transmission capacity again provides system value.
Figure 3: Comparison of system (left) and flexible consumer costs (right) without load shifting (blue) and with decentralized load shifting (orange) for a range of transmission capacities.
V Conclusion
In this paper, we develop an analytical framework for understanding how price-responsive flexible consumers interact with power system operations through marginal price signals. By deriving closed-form expressions for both private and social objectives, we characterize when decentralized load-shifting behavior aligns with, or diverges from, system-optimal dispatch. We further analyze the marginal value of transmission capacity in the presence of flexible load and show that conventional investment signals—such as transmission shadow prices—may fail to reflect the true cost-reducing value of capacity expansion when flexible consumers adjust their demand endogenously. This highlights the need to account for strategic load response in capacity expansion modeling as well as designing pricing mechanisms and coordination policies for future power systems with large flexible loads.
Acknowledgments
The authors would like to thank Nathan Engelman Lado, Rahman Khorramfar, and Daniel Shen for their insights as well as the MIT Energy Initiative and MIT Climate Grand Challenges, which contributed to the development of this work.